MICROWAVE SPECTROSCOPY OF TRANS-ETHYL METHYL ETHER IN THE TORSIONALLY EX-CITED STATE 3

\underline{KAORI KOBAYASHI}, KEIGO MURATA, SHOZO TSUNEKAWA, Department of Physics, University of Toyama, 3190 Gofuku, Toyama, 930-8555 Japan; and NOBUKIMI OHASHI, Kanazawa University, Japan.

The trans-ethyl methyl ether molecule (CH_{3}CH_{2}OCH_{3}) has two methyl group internal rotors which are equivalent to the two vibrational motions, ν_{28} and ν_{29}. There is another low-lying torsional motion which is a skeltal torsion (ν_{30}) and does not cause splitting. The microwave spectra of the trans-ethyl methyl ether molecule in the ν_{28} = 1, ν_{29} = 1, and ν_{30} = 1 have been studied and interactions between these states were discussed. In this paper we report results on the ν_{30} = 2, and 3 state. The analysis based on Hougen’s tunneling matrix formulation considering two methyl groups are used. We try to interpret tunneling parameters obtained in the present analysis quantitatively from the viewpoint of torsion-torsion interaction.